Differentiate with respect to x , x³-2x from the first principle


1
Expert's answer
2021-04-27T13:46:56-0400

(i) f(x)=x

By first principle-

"f'(x)=lim_{h\rightarrow 0}\dfrac{f(x+h)-f(x)}{h}=lim_{h\rightarrow 0}\dfrac{x+h-x}{h}=lim_{h\rightarrow 0}\dfrac{h}{h}=1"


(ii) "f(x)=x^3-2x"


By first principle-


 "f'(x)=lim_{h\rightarrow 0}\dfrac{f(x+h)-f(x)}{h}"


"=lim_{h\rightarrow 0}\dfrac{(x+h)^3-2(x+h)-(x^3-2x)}{h}"


"=lim_{h\rightarrow 0}\dfrac{(x^3+h^3+3x^2h+3xh^2-2x-2h-x^3-2x)}{h}"


"=lim_{h\rightarrow 0}\dfrac{(h^3+3x^2h+3xh^2-2h)}{h}"


"=lim_{h\rightarrow 0}\dfrac{h(h^2+3x^2+3xh-2)}{h}"


"=lim_{h\rightarrow 0}(h^2+3x^2+3xh-2)=3x^2-2"


Hence "f'(x)=3x^2-2"